OpenStudy (anonymous):

Bert is planning to open a savings account that earns 1.6% simple interest yearly. He wants to earn exactly $192 in interest after 2 years. How much money should he deposit?

OpenStudy (anonymous):

will give a medal!!!

OpenStudy (pixiedust1):

Okay! The formula for simple interest is Investment Amount*Time*Percent=Interest n can represent the amount of the original investment (since we do not know it) So, its n*2*0.016=192 Now, let's do some guess and check for the value of n. :D (I did this work elsewhere, so I'll show you my answer after guessing and checking.) Let's try $6000 6000*2*0.016=$192 Therefore, Bert will have to invert $6000 to get $192 simple interest after 2 years at 1.6%

OpenStudy (anonymous):

thanks and also 2. Mr. Smith lives in State Line City on the border of two states. He buys a CD player for $86.29 in the state with 6% sales tax. He then crosses the state line and buys 7 CDs at $15.49 each in the state that has 7% sales tax. How much does he spend in all?

OpenStudy (anonymous):

my chios are$108.04 $207.49 $91.47 $199.90

OpenStudy (pixiedust1):

Okay! Here goes Let's figure out for the $86.29 in the state with 6% sales tax. To find the tax, we multiply 86.39*0.06=

OpenStudy (pixiedust1):

What does that equal?

OpenStudy (anonymous):

5.1774

OpenStudy (pixiedust1):

I have to go soon, so I'm just gonna finish this up for you. 86.39*0.06=5.17 in tax Now, we add the tax to the purchase price. 86.39*5.17=$91.46 for the cd player in the state with 6% sales tax Now, let's do the other state. Let's see how much he's paying for 1 cd with tax. 15.49*0.07=$1.08 in tax per cd Now, we add the tax amount in for the 1 cd. 15.49+1.08=$16.57 per 1 cd Now, let's multiply the amount per cd plus tax by the 7 cds 16.57*7=$115.99 for the 7 cds Now, we know the player was 91.46 and the CD was 115.99 So, 115.99+91.46=$207.99 for everything

OpenStudy (anonymous):

The price of a jacket is initially marked at $39.99. The sign above the rack says the jacket is 20% off. Sales tax is 7.25%. What is the total cost of the jacket?

OpenStudy (pixiedust1):

39.99*0.20=7.99 39.99-7.99=32 32*0.0725=$2.32 32+2.32=$34.32

OpenStudy (pixiedust1):

@loltk5534